OpenStudy (suwhitney):

Find the inverse of f(x)=-5x+4/7-6x

OpenStudy (suwhitney):

I understand that you first make the f(x) a Y. and then you swith all of the X's and Y's, ending up with x=-5y+4/7-6y, but after that I am stuck.

OpenStudy (anonymous):

first you want to combine like terms to make it easier to find the inverse: f(x) = -11x + 4/7 y = -11x + 4/7 write in terms of x x = -11y+ 4/7 and solve for y: x-4/7 = -11y divide both sides by -11 (x-4/7)/-11 = y x/-11 + 4/-77 = y

OpenStudy (suwhitney):

Hmm.. not sure if I understand. Because I know that the right answer is f^-1=7x-4/6x-5. Im just not sure of the exact steps taken...

OpenStudy (anonymous):

i would combine like terms first, i don't understand why you wouldn't. maybe i'm thinking about it wrong. sorry. i tried

OpenStudy (suwhitney):

It's okay.. I missplaced by scratch paper where I wrote down how he worked it out but I think it had something to do with multiplying top and bottom by the negative recripicol. (sp?) and then going from there.. thanks for trying :)